OpenStudy (anonymous):

Can someone help me with solve 7m-(2m-9)=39

OpenStudy (anonymous):

hey, how's it going?

OpenStudy (anonymous):

hi can you help me out with this equation?

OpenStudy (anonymous):

sure, so by the distributive property, 7m-2m-(-9) = 39

OpenStudy (anonymous):

5m+9=39, subtract 9 from both sides, 5m+9-9=39-9, 5m = 30

OpenStudy (anonymous):

divide both sides by 5, (5m)/5 = 30/5 , m = 6